# Hybrid inflation and gravitational waves from accidentally light scalars

### v1

@@ 3 Inflation parameters and CMB observables
As inflation proceeds along the tree-level flat direction, the waterfall field $\chi$ is stabilized at $\chi=0$ by its effective mass, hence the model qualifies as a single-field model. The evolution of scalar and tensor perturbations is then dictated by the potential in Eq. (6), with the slow-roll parameters defined by
@@
At the leading order in the slow-roll parameters, the tensor-to-scalar ratio $r$, the amplitude of the curvature power spectrum $A_{s}$, the spectral index of the curvature power spectrum $n_{s}$, the running of the spectral index $n_{r}$ and the running of the running $n_{rr}$ can be computed as
@@
The independent model parameters entering the calculation of the CMB power spectrum can be taken as $V_{0}$, $M$, $f$ and $a_{*}$. Considering that $V_{0}\gg M^{4}$, the observables are
@@
We identify the field value $a_{\rm end}$ where inflation ends with the critical point $a_{c}=6f\alpha_{c}$, where the waterfall field becomes tachyonic, see Eq. (9). We will comment on the assumptions underlying this identification at the end of this section. As the end of inflation is controlled by the coupling between the inflaton and the waterfall field, $a_{\rm end}$ is an independent parameter of our model, which can be freely adjusted: by suitably choosing $a_{\rm end}$, we are always able to accommodate enough $e$-folds of inflation. For concreteness, in what follows we fix $N=60$ as a reference value.
A characteristic feature of the model is to predict positive $n_{r}$ and $n_{rr}$. A second feature concerns the running of the running of the spectral index, which for this model is not independent from $n_{s}$ and $n_{r}$, $n_{rr}\simeq(1-n_{s})n_{r}$, leading to the constraint $n_{rr}<4\times 10^{-4}$. This is more stringent than the direct experimental bound on $n_{rr}$. The upcoming satellite mission SPHEREx [26] will measure the running of the spectral index with a target sensitivity $\Delta n_{r}\sim 10^{-3}$, and therefore it could rule out the model in the near future.
The four parameters of the model are not all determined by the CMB observables. Requiring that the model parameters do not suffer from fine-tuning problems will lead to further relations between them, see section 3.1. To better understand how observation constrains them, we invert the relations given in Eq. (LABEL:eq:CMBObs) and obtain
@@
By imposing $A_{s}=2.105\times 10^{-9}$, $n_{s}=0.965$, $n_{r}<0.011$ and $r<0.036$, we can derive the inequalities
@@
In the left-hand panel of Fig. 2 we chart the region allowed by observations in the $(n_{r},r)$ plane, once $A_{s}$ and $n_{s}$ are fixed. All the above relations have been derived under the assumption that $V_{0}$ drives inflation, i.e. $V_{0}\gg M^{4}$. Below we quantify what separation is needed between the two scales, in order not to end inflation before the critical point $a=a_{c}$. In particular, we require $\epsilon<1$ all along the inflationary trajectory. [^4] For slow-roll dynamics to persists we also have to check that $\lvert\eta\rvert<1$ for any value of $a$, but this leads to a constraint which is very similar to Eq. (19).
@@
$\frac{M^{4}}{V_{0}}<\sqrt{\frac{2f^{2}}{2f^{2}+M_{\mathrm{P}}^{2}}}\iff r<n_{r}\left(-1+\sqrt{1+\frac{64}{2n_{r}+(n_{s}-1)^{2}}}\right)\,.$ (19)
The region where this condition is not satisfied is shaded in dark grey in the left-hand panel of Fig. 2.
Figure 2: Left-hand panel. The values of $M$ and $f$ in units of $M_{\mathrm{P}}$, as a function of $n_{r}$ and $r$, while fixing $A_{s}=2.105\times 10^{-9}$ and $n_{s}=0.965$. The largest displayed values of $n_{r}$ and $r$ correspond to the present experimental upper bounds. Solid (dashed) lines are contours of constant $M$ ($f$). The dark-grey region violates the slow-roll condition before the critical point. Right-hand panel. The white region is the one which complies with naturalness criteria, according to the discussion in section  3.1. In the orange regions the value of $f$ is unnaturally small, while in the red region $f>M_{\mathrm{P}}$. The parameter space covered by the blue and green regions requires a significant fine-tuning of the inflation initial and final conditions, respectively.
In Fig. 3 we show the predictions of the model in the $(n_{s},r)$ plane, together with the region allowed by the latest CMB measurements [2, 25]. Our model is in agreement with observations for natural values of $f$, i.e. for $f$ smaller than, but not too far from $M_{\mathrm{P}}$.
Figure 3: The purple areas show the 95% CL and 68 % CL preferred regions from Planck 2018 + BAO + lensing + BICEP/Keck 2018 data [25]. We fixed the amplitude of the scalar power spectrum at $A_{s}=2.105\times 10^{-9}$. In the left-hand panel we set the running of the spectral index at $n_{r}=10^{-3}$, while in the right-hand panel we took $n_{r}=10^{-4}$. The horizontal, green lines are obtained by fixing $f=0.2\,M_{\mathrm{P}}$ and varying $M$, while along the vertical, blue lines $M/f=3\times 10^{-3}$ is fixed and $f$ varies. The orange and the red regions are disfavoured by naturalness (see section 3.1): low values of the tensor-to-scalar ratio $r$ require unnaturally small $f$ while a large $r$ is associated to transplanckian $f$. The relation between $f$ and $r$ depends on the value assumed for $n_{r}$, according to Eq. (16).
We require the end of inflation to occur within one Hubble time after the critical point has been reached by the accident. This allows us to identify $a_{c}=6f\alpha_{c}$ with the point $a_{\rm end}$ in field space in which inflation ends. For this requirement to hold, the Hubble rate at the critical point, $H_{c}$, must be much lower than the value of the waterfall mass, after $\Delta t\sim H_{c}^{-1}$ has elapsed since $a=a_{c}$
@@
where $\Delta a$ is the accident excursion during $\Delta t$. This condition is also known as the “waterfall condition” [7]. Given that the energy density during inflation is dominated by the false vacuum, the Hubble rate can be approximated to $H_{c}\simeq V_{0}/3M_{\mathrm{P}}^{2}$. Moreover, the evolution of $a$ is governed, in the slow-roll approximation, by $3\,H_{c}\,\dot{a}\simeq\partial V_{\rm eff}/\partial a$. With these considerations in mind the above waterfall condition translates into a lower bound on the inflaton-waterfall quartic coupling,
@@
By making use of the relations in Eqs. (16) and (15), we find that the right-hand side of the above condition is a monotonic, increasing function of $n_{r}$, once $A_{s}$, $n_{s}$ and $N$ are fixed. For $n_{r}$ taking its largest value allowed by observations, Eq. (21) becomes $\zeta\gg 10^{-6}$, meaning that we can always choose a perturbative value of $\zeta$ for which the waterfall condition holds. We note that the requirement for inflation to end as soon the critical point is reached is not imposed by observations. In principle, one could also allow inflation to proceed along the waterfall direction for some 20 to 40 $e$-folds as realised in the so called “mild hybrid inflation” scenario [27, 28, 29].
@@ 3.1 Naturalness considerations
The aim of this section is to asses which portion of the parameter space allowed by observations complies with naturalness. An important motivation for considering this model in the first place is that the inflaton potential is nearly flat for generic values of the couplings, without requiring fine-tuning. To what extent does the model remain “natural”, or minimally fine-tuned, once we impose that it should give rise to a quantitatively realistic phenomenology?
It turns out that there is a preferred range of natural values for the running of the spectral index $n_{r}$. A lower bound on $n_{r}$ is obtained as follows: from Eq. (LABEL:eq:CMBObs) we have
@@
one obtains the naturalness bound $n_{r}\gtrsim 6.5\times 10^{-5}$. Requiring the fine-tuning to be less than 1% gives, instead, $n_{r}\gtrsim 6\times 10^{-7}$. Such lower bounds are shown as blue-shaded regions in the right-hand panel of Fig. 2.
An upper bound on $n_{r}$ is similarly given by the following argument: fixing the number $N$ of observable e-folds to $N=60$, the parameter combination $f^{2}V_{0}/(M_{\mathrm{P}}^{2}M^{4})$ could in principle be made arbitrarily small by tuning the ending point of inflation $a_{\rm end}$ close to zero, see Eq. (15) (and by simultaneously tuning $a_{*}$ close to $a_{*}=\frac{\pi}{2}f$ to obtain the observed value of $n_{s}$ according to Eqs. (LABEL:eq:CMBObs)). This would imply, again by Eqs. (LABEL:eq:CMBObs), that $n_{r}$ could be arbitrarily large. But $a_{\rm end}$ is given by a combination of independent model parameters which have no reason to conspire to produce $a_{\rm end}\simeq 0$. Imposing for concreteness
@@
this yields $n_{r}\lesssim 2\times 10^{-3}$, shown by green shading in the right-hand panel of Fig. 2.
@@
Since $f^{2}=\mu_{\phi}^{2}/6\lambda_{\phi}$, this implies $f\gtrsim 0.1\,M_{\mathrm{P}}$, under the technically natural assumption that $\varepsilon$ and $\zeta$ are at most of the order of $\lambda_{\phi}$. Therefore, the orange-shaded region in the right-hand panel of Fig. 2 is disfavoured. From Eqs. (LABEL:eq:CMBObs) it follows that $r=4\,n_{r}\,(f/M_{\mathrm{P}})^{2}$ and thus
@@
shown in orange shading in Fig. 3. Depending on how much fine-tuning one is willing to accept, this bound can of course shift by several orders of magnitude, as also indicated in the right-hand panel of Fig. 2.
On the other hand, values of $f$ too close to $\Lambda_{\rm UV}$ imply that higher-dimensional operators, which may spoil the tree-level flatness of the inflaton potential, can no longer be neglected. In that sense, the “least fine-tuned” region of parameter space is also the one with the least theoretical control. In Figs. 2 and 3 we shaded in red the regions with $f\gtrsim M_{\mathrm{P}}$, or equivalently $r\gtrsim 4n_{r}$.
Such naturalness constraints translate into some bounds on the model parameters. For instance, the inflationary scale is given by $V_{0}\simeq\lambda_{\chi}v_{\chi}^{4}/4$, and the above bounds on $r$ and $n_{r}$ can be translated into bounds on $V_{0}$ once $A_{s}$ is fixed, according to Eq. (16). This, together with $0.1\,M_{\mathrm{P}}\lesssim v_{\chi}\lesssim M_{\mathrm{P}}$ (we require $v_{\chi}$ to be subplanckian, analogously to $f$) gives a natural range of values for the $\chi$ quartic coupling,
@@ 4 Gravitational waves from preheating
Hybrid inflation is followed by a stage of tachyonic preheating [8, 9]. The tachyonic instability of the waterfall field exponentially enhances field inhomogeneities. Such inhomogeneous, bubble-like structures behave as scalar waves and scatter off each other until the Universe is homogenized and driven to local thermal equilibrium. A stochastic gravitational wave background (SGWB) is produced in the process [30, 31, 32, 33]. GWs quickly decouple from the background dynamics and propagate (almost) undisturbed throughout cosmic history, hence carrying information about the process that sourced them. Given the energy scale of inflation, which is typically large, the frequency at which GWs are produced during preheating lies well beyond the reach of current or even planned interferometers. However, it has been shown, e.g. in Ref. [33], that by sufficiently lowering the scale of inflation, the signal might be brought within the projected sensitivity of the Einstein Telescope (ET) [34]. It is also worth mentioning that there have been proposals of using optical atomic clocks to detect high-frequency GWs, see e.g. Ref. [35] and references therein.
The spectrum of GWs produced during preheating presents a peaked shape, with the frequency and amplitude at the peak given by [33]
@@
where $k_{\star}$ is the typical momentum enhanced by the tachyonic instability, related to the characteristic size of inhomogeneities as $R_{\star}\sim 1/k_{\star}$. [^5] Here $k_{\star}$ is unrelated to the pivot scale of CMB observables $k_{*}$ of Section 3. Here, $\rho_{c}\simeq V_{0}$ and $H_{c}^{2}\simeq V_{0}/(3M_{\mathrm{P}}^{2})$ are the energy density and the Hubble rate at the end of inflation or, analogously, at the onset of the tachyonic instability. The value of $k_{\star}$ depends on the inflationary dynamics close to the waterfall transition.
From Eq. (28) one may already deduce that for a signal of GWs from preheating to be within the reach of the ET, the inflation scale must lie in the range
@@
which is associated to an unobservably small tensor-to-scalar ratio, according to Eq. (16). Should $r$ be measured in the future, along with a SGWB at the ET which can be traced back to tachyonic preheating, this would therefore rule out our model (and other models of hybrid inflation, where tachyonic preheating is an unavoidable feature).
In principle, two different regimes for GW production from preheating can be identified [33], depending on the velocity $\dot{a}_{c}$ of the inflaton at the critical point $a=a_{c}$ where the waterfall field becomes tachyonic. In the first regime, $a$ is fast enough for the tachyonic instability to be driven by its classical dynamics. Another possibility is that $\dot{a}_{c}$ is sufficiently small, so that preheating is triggered by quantum fluctuations. In the present work, we take model parameters such that inflation ends as soon as the inflaton crosses the critical point, see Eq. (21). In that case classical rolling always dominates.
We are interested in an order-of-magnitude estimate of the amplitude and frequency of the GWs signal at the peak. Due to the fact that tachyonic preheating is a non-linear process, a more precise computation of the GW spectrum would require lattice simulations which is beyond the scope of this work. In the slow-roll approximation, $\ddot{a}\ll 3H\dot{a}$, the inflaton velocity at the critical point reads
@@
Figure 4: Sensitivity region (yellow) of the ET [34] for the GW spectrum produced during tachyonic preheating, as a function of $n_{r}$ and $r$ (or equivalently $V_{0}$). In the left-hand panel we take $\tilde{\mu}_{\chi}^{2}/f^{2}=1$, while in the right-hand one we fix $\tilde{\mu}_{\chi}^{2}/f^{2}=10^{-6}$. The projected sensitivity for the ET is the one derived in Ref. [11] and it is shown in Fig. 5. To guide the eye we also plot lines of constant peak frequency and amplitude for the preheating GW spectrum. The brown region shows the integrated sensitivity for the third observing run of the LIGO-Virgo-KAGRA network [10].

### v2

@@ 3 Inflation parameters and CMB observables
As inflation proceeds along the tree-level flat direction, the waterfall field $\chi$ is stabilized at $\chi=0$ by its effective mass; hence, the model qualifies as a single-field model. The evolution of scalar and tensor perturbations is then dictated by the potential in Eq. (6), with the slow-roll parameters defined by
@@
At the leading order in the slow-roll parameters, the tensor-to-scalar ratio $r$, the amplitude of the curvature power spectrum $A_{s}$, the spectral index of the curvature power spectrum $n_{s}$, the running of the spectral index $n_{r}$, and the running of the running $n_{rr}$ can be computed, respectively, as
@@
The independent model parameters entering the calculation of the CMB power spectrum can be taken as $V_{0}$, $M$, $f$, and $a_{*}$. Considering that $V_{0}\gg M^{4}$, the observables are
@@
We identify the field value $a_{\rm end}$ where inflation ends with the critical point $a_{c}=6f\alpha_{c}$, where the waterfall field becomes tachyonic; see Eq. (9). We will comment on the assumptions underlying this identification at the end of this section. As the end of inflation is controlled by the coupling between the inflaton and the waterfall field, $a_{\rm end}$ is an independent parameter of our model, which can be freely adjusted: By suitably choosing $a_{\rm end}$, we are always able to accommodate enough $e$-folds of inflation. For concreteness, in what follows, we fix $N=60$ as a reference value.
A characteristic feature of the model is to predict positive $n_{r}$ and $n_{rr}$. A second feature concerns the running of the running of the spectral index, which for this model is not independent from $n_{s}$ and $n_{r}$, $n_{rr}\simeq(1-n_{s})n_{r}$, leading to the constraint $n_{rr}<4\times 10^{-4}$. This is more stringent than the direct experimental bound on $n_{rr}$. The upcoming satellite mission SPHEREx [30] will measure the running of the spectral index with a target sensitivity $\Delta n_{r}\sim 10^{-3}$, and, therefore, it could rule out the model in the near future.
The four parameters of the model are not all determined by the CMB observables. Requiring that the model parameters do not suffer from fine-tuning problems will lead to further relations between them; see Sec. 3.1. To better understand how observation constrains them, we invert the relations given in Eq. (LABEL:eq:CMBObs) and obtain
@@
By imposing $A_{s}=2.105\times 10^{-9}$, $n_{s}=0.965$, $n_{r}<0.011$, and $r<0.036$, we can derive the inequalities
@@
In the left-hand panel in Fig. 2, we chart the region allowed by observations in the $(n_{r},\,r)$ plane, once $A_{s}$ and $n_{s}$ are fixed. All the above relations have been derived under the assumption that $V_{0}$ drives inflation, i.e., $V_{0}\gg M^{4}$. Below, we quantify what separation is needed between the two scales, in order not to end inflation before the critical point $a=a_{c}$. In particular, we require $\epsilon<1$ all along the inflationary trajectory. [^4] For slow-roll dynamics to persist, we also have to check that $\lvert\eta\rvert<1$ for any value of $a$, but this leads to a constraint which is very similar to Eq. (20).
@@
$\frac{M^{4}}{V_{0}}<\sqrt{\frac{2f^{2}}{2f^{2}+M_{\mathrm{P}}^{2}}}\,,$ (19)
or, equivalently,
$r<n_{r}\left(-1+\sqrt{1+\frac{64}{2n_{r}+(n_{s}-1)^{2}}}\right)\,.$ (20)
The region where this condition is not satisfied is shaded in dark gray in the left-hand panel in Fig. 2.
Figure 2: Left-hand panel. The values of $M$ and $f$ in units of $M_{\mathrm{P}}$, as a function of $n_{r}$ and $r$, while fixing $A_{s}=2.105\times 10^{-9}$ and $n_{s}=0.965$. The largest displayed values of $n_{r}$ and $r$ correspond to the present experimental upper bounds. Solid (dashed) lines are contours of constant $M$ ($f$). The dark-gray region violates the slow-roll condition before the critical point. Right-hand panel. The white region is the one which complies with naturalness criteria, according to the discussion in Sec.  3.1. In the orange regions the value of $f$ is unnaturally small, while in the red region $f>M_{\mathrm{P}}$. The parameter space covered by the blue and green regions requires a significant fine-tuning of the inflation initial and final conditions, respectively.
In Fig. 3, we show the predictions of the model in the $(n_{s},\,r)$ plane, together with the region allowed by the latest CMB measurements [2, 29]. Our model is in agreement with observations for natural values of $f$, i.e., for $f$ smaller than, but not too far from $M_{\mathrm{P}}$.
Figure 3: The purple areas show the 95% CL and 68 % CL preferred regions from Planck 2018 + BAO + lensing + BICEP/Keck 2018 data [29]. We fixed the amplitude of the scalar power spectrum at $A_{s}=2.105\times 10^{-9}$. In the left-hand panel we set the running of the spectral index at $n_{r}=10^{-3}$, while in the right-hand panel we took $n_{r}=10^{-4}$. The horizontal, green lines are obtained by fixing $f=0.2\,M_{\mathrm{P}}$ and varying $M$, while along the vertical, blue lines $M/f=3\times 10^{-3}$ is fixed and $f$ varies. The orange and the red regions are disfavoured by naturalness (see section 3.1): Low values of the tensor-to-scalar ratio $r$ require unnaturally small $f$, while a large $r$ is associated to transplanckian $f$. The relation between $f$ and $r$ depends on the value assumed for $n_{r}$, according to Eq. (16).
We require the end of inflation to occur within one Hubble time after the critical point has been reached by the accident. This allows us to identify $a_{c}=6f\alpha_{c}$ with the point $a_{\rm end}$ in field space in which inflation ends. For this requirement to hold, the Hubble rate at the critical point, $H_{c}$, must be much lower than the value of the waterfall mass, after $\Delta t\sim H_{c}^{-1}$ has elapsed since $a=a_{c}$:
@@
where $\Delta a$ is the accident excursion during $\Delta t$. This condition is also known as the “waterfall condition” [7]. Given that the energy density during inflation is dominated by the false vacuum, the Hubble rate can be approximated to $H_{c}\simeq V_{0}/3M_{\mathrm{P}}^{2}$. Moreover, the evolution of $a$ is governed, in the slow-roll approximation, by $3\,H_{c}\,\dot{a}\simeq\partial V_{\rm eff}/\partial a$. With these considerations in mind, the above waterfall condition translates into a lower bound on the inflaton-waterfall quartic coupling:
@@
By making use of the relations in Eqs. (16) and (15), we find that the right-hand side of the above condition is a monotonic, increasing function of $n_{r}$, once $A_{s}$, $n_{s}$, and $N$ are fixed. For $n_{r}$ taking its largest value allowed by observations, Eq. (22) becomes $\zeta\gg 10^{-6}$, meaning that we can always choose a perturbative value of $\zeta$ for which the waterfall condition holds. We note that the requirement for inflation to end as soon the critical point is reached is not imposed by observations. In principle, one could also allow inflation to proceed along the waterfall direction for some 20 -- 40 $e$-folds as realized in the so called “mild hybrid inflation” scenario [31, 32, 33].
@@ 3.1 Naturalness considerations
The aim of this section is to assess which portion of the parameter space allowed by observations complies with naturalness. An important motivation for considering this model in the first place is that the inflaton potential is nearly flat for generic values of the couplings, without requiring fine-tuning. To what extent does the model remain “natural”, or minimally fine-tuned, once we impose that it should give rise to a quantitatively realistic phenomenology?
It turns out that there is a preferred range of natural values for the running of the spectral index $n_{r}$. A lower bound on $n_{r}$ is obtained as follows: From Eq. (LABEL:eq:CMBObs), we have
@@
one obtains the naturalness bound $n_{r}\gtrsim 6.5\times 10^{-5}$. Requiring the fine-tuning to be less than 1% gives, instead, $n_{r}\gtrsim 6\times 10^{-7}$. Such lower bounds are shown as blue-shaded regions in the right-hand panel in Fig. 2.
An upper bound on $n_{r}$ is similarly given by the following argument: Fixing the number $N$ of observable e-folds to $N=60$, the parameter combination $f^{2}V_{0}/(M_{\mathrm{P}}^{2}M^{4})$ could, in principle, be made arbitrarily small by tuning the ending point of inflation $a_{\rm end}$ close to zero; see Eq. (15) (and by simultaneously tuning $a_{*}$ close to $a_{*}=\frac{\pi}{2}f$ to obtain the observed value of $n_{s}$ according to Eqs. (LABEL:eq:CMBObs)). This would imply, again by Eqs. (LABEL:eq:CMBObs), that $n_{r}$ could be arbitrarily large. But $a_{\rm end}$ is given by a combination of independent model parameters which have no reason to conspire to produce $a_{\rm end}\simeq 0$. Imposing for concreteness
@@
this yields $n_{r}\lesssim 2\times 10^{-3}$, shown by green shading in the right-hand panel in Fig. 2.
@@
Since $f^{2}=\mu_{\phi}^{2}/6\lambda_{\phi}$, this implies $f\gtrsim 0.1\,M_{\mathrm{P}}$, under the technically natural assumption that $\varepsilon$ and $\zeta$ are at most of the order of $\lambda_{\phi}$. Therefore, the orange-shaded region in the right-hand panel in Fig. 2 is disfavoured. From Eqs. (LABEL:eq:CMBObs), it follows that $r=4\,n_{r}\,(f/M_{\mathrm{P}})^{2}$ and, thus,
@@
shown in orange shading in Fig. 3. Depending on how much fine-tuning one is willing to accept, this bound can, of course, shift by several orders of magnitude, as also indicated in the right-hand panel in Fig. 2.
On the other hand, values of $f$ too close to $\Lambda_{\rm UV}$ imply that higher-dimensional operators, which may spoil the tree-level flatness of the inflaton potential, can no longer be neglected. In that sense, the “least fine-tuned” region of parameter space is also the one with the least theoretical control. In Figs. 2 and 3, we shaded in red the regions with $f\gtrsim M_{\mathrm{P}}$, or, equivalently, $r\gtrsim 4n_{r}$.
Such naturalness constraints translate into some bounds on the model parameters. For instance, the inflationary scale is given by $V_{0}\simeq\lambda_{\chi}v_{\chi}^{4}/4$, and the above bounds on $r$ and $n_{r}$ can be translated into bounds on $V_{0}$ once $A_{s}$ is fixed, according to Eq. (16). This, together with $0.1\,M_{\mathrm{P}}\lesssim v_{\chi}\lesssim M_{\mathrm{P}}$ (we require $v_{\chi}$ to be subplanckian, analogously to $f$), gives a natural range of values for the $\chi$ quartic coupling:
@@ 4 Gravitational waves from preheating
Hybrid inflation is followed by a stage of tachyonic preheating [8, 9]. The tachyonic instability of the waterfall field exponentially enhances field inhomogeneities. Such inhomogeneous, bubble-like structures behave as scalar waves and scatter off each other until the Universe is homogenized and driven to local thermal equilibrium. A stochastic gravitational wave background (SGWB) is produced in the process [34, 35, 36, 37]. GWs quickly decouple from the background dynamics and propagate (almost) undisturbed throughout cosmic history, hence carrying information about the process that sourced them. Given the energy scale of inflation, which is typically large, the frequency at which GWs are produced during preheating lies well beyond the reach of current or even planned interferometers. However, it has been shown, e.g., in Ref. [37], that, by sufficiently lowering the scale of inflation, the signal might be brought within the projected sensitivity of the Einstein Telescope (ET) [38]. It is also worth mentioning that there have been proposals of using optical atomic clocks to detect high-frequency GWs; see, e.g., Ref. [39] and references therein.
The spectrum of GWs produced during preheating presents a peaked shape, with the frequency and amplitude at the peak given, respectively, by [37]
@@
where $k_{\star}$ is the typical momentum enhanced by the tachyonic instability, related to the characteristic size of inhomogeneities as $R_{\star}\sim 1/k_{\star}$. [^5] Here $k_{\star}$ is unrelated to the pivot scale of CMB observables $k_{*}$ in Sec. 3. Here, $\rho_{c}\simeq V_{0}$ and $H_{c}^{2}\simeq V_{0}/(3M_{\mathrm{P}}^{2})$ are, respectively, the energy density and the Hubble rate at the end of inflation or, analogously, at the onset of the tachyonic instability. The value of $k_{\star}$ depends on the inflationary dynamics close to the waterfall transition.
From Eq. (29), one may already deduce that, for a signal of GWs from preheating to be within the reach of the ET, the inflation scale must lie in the range
@@
which is associated to an unobservably small tensor-to-scalar ratio, according to Eq. (16). Should $r$ be measured in the future, along with a SGWB at the ET which can be traced back to tachyonic preheating, this would, therefore, rule out our model (and other models of hybrid inflation, where tachyonic preheating is an unavoidable feature).
In principle, two different regimes for GW production from preheating can be identified [37], depending on the velocity $\dot{a}_{c}$ of the inflaton at the critical point $a=a_{c}$ where the waterfall field becomes tachyonic. In the first regime, $a$ is fast enough for the tachyonic instability to be driven by its classical dynamics. Another possibility is that $\dot{a}_{c}$ is sufficiently small, so that preheating is triggered by quantum fluctuations. In the present work, we take model parameters such that inflation ends as soon as the inflaton crosses the critical point; see Eq. (22). In that case, classical rolling always dominates.
We are interested in an order-of-magnitude estimate of the amplitude and frequency of the GWs signal at the peak. Because of the fact that tachyonic preheating is a non-linear process, a more precise computation of the GW spectrum would require lattice simulations which are beyond the scope of this work. In the slow-roll approximation, $\ddot{a}\ll 3H\dot{a}$, the inflaton velocity at the critical point reads
@@
Figure 4: Sensitivity region (yellow) of the ET [38] for the GW spectrum produced during tachyonic preheating, as a function of $n_{r}$ and $r$ (or equivalently $V_{0}$). In the left-hand panel we take $\tilde{\mu}_{\chi}^{2}/f^{2}=1$, while in the right-hand panel we fix $\tilde{\mu}_{\chi}^{2}/f^{2}=10^{-6}$. The projected sensitivity for the ET is the one derived in Ref. [15], and it is shown in Fig. 5. To guide the eye, we also plot lines of constant peak frequency and amplitude for the preheating GW spectrum. The brown region shows the integrated sensitivity for the third observing run of the LIGO-Virgo-KAGRA network [14].